RC meeting and Run Plan, Tuesday December 13th
From SBS wiki
Return to the RC Meeting Calendar
Contents
RC Meeting Information
Current RC: | Eric King [ericking (@jlab.org or @temple.edu)] | |
Time: | Daily at 3:30PM | |
Location: | Counting House Conference Room |
Remote Connection Details
- Join ZoomGov Meeting
- https://jlab-org.zoomgov.com/j/1619032184?pwd=aWhTVWxya1ZZQzA1K2ZPSTd2ZX...
- Meeting ID: 161 903 2184
- Passcode: 260715
- Dial by your location
- +1 669 254 5252 US (San Jose)
- +1 646 828 7666 US (New York)
- +1 551 285 1373 US
- +1 669 216 1590 US (San Jose)
- 833 568 8864 US Toll-free
- Find your local number: https://jlab-org.zoomgov.com/u/acj3pTicMx
Shift Schedule
Until Shutdown
All shifts are full! :)
Accelerator Status
Lost some production time due to accelerator issues
- Tuesday OWL -- 2L12 dropped out of HV ~3h15m [DT Event ID: 17476]
- Tuesday OWL -- Bypass 2L21 ~30m [DT Event ID: 17475]
- Tuesday OWL -- 2L21 dropped out of HV ~20m [DT Event ID: 17474]
Current Status
Took production data on 3-He during OWL, Moller measurement and SBS target systematic studies during DAY (Moller HH magnet quenched), return to production or Moller during SWING.
Shift Summaries
Things That Happened
On the Hall side things were largely uneventful during SWING and OWL.
- Moller magnet quenched at 62 amps, measurement cancelled. 4100773
- Moller magnet cooled back down. 4100805
- Don and I spoke with Bogdan and came to the agreement to not turn production beam back on yet and try to re-train the Moller magnet.
- BB HCal GUI froze while turning off Detector HV this morning; Jiwan turned off using secondary method.
- Difficulty turning back on.
- Sebastian and Bob came in to help troubleshoot. Hall access was made and HV card replaced. 4100819
Beam Current
We've been happily running at 40uA when taking production data.
Target
Target polarization has been stable. List of the recent polarizations
- TUE 14:20 → 43.9% (Beam off for a while) 4100829
- TUE 08:00 → 40.8% (Beam off 2hr) 4100714
- TUE 07:09 → 41.6% (Beam off 1hr) 4100705
- TUE 06:05 → 42.8% 4100691
- TUE 01:00 → 41.2% 4100601
- MON 20:50 → 41.6% 4100534
- MON 17:10 → 42.0% 4100486
Møller Measurement
Moller target magnet quenched. Perhaps re-do measurement during SWING?
Issues remaining
- BB GEM APV unreachable and cannot be repaired at this time.
Short Term Run Plan
- Continue production run on He3 target, 40 uA
- Use GEnII-noSBSGEMs DAQ config
- Use ps3=0, ps5=6, others =-1
- BBCal threshold at -366 mV, HCal threshold at -19 mV
- NMR measurement every 3 - 4 hours, close to 4 hours is better
Planned Accesses to the Hall
None
Special Requests
Please email me ericking (@jlab.org or @temple.edu) if there's something to be added.
Notices in Effect for Shift Crew
- Power down SBS, BB, and Detector HV before making an access to the Hall
Data Collection Status
- 78 out of 141 days (~55.3%)
- Total uncorrected accumulated charge on He3: 57.710 C out of 232C (24.86%)
- 2-Pass: 13.467 C
- 3-Pass: 44.243 C (Tuesday 8AM)